# Complementarity between Cosmic String Gravitational Waves and long lived particle searches in laboratory

### v1

@@ VI Complementarity between GW and laboratory searches
Probing the aforementioned long-lived scalar $S$ in upcoming colliders is very challenging. However, in the early universe, non-negligible processes such as inflaton decay or moduli decay can result in a significant initial abundance of $S$, potentially altering the standard scenario. For freeze-out, a sufficiently high interaction rate would eliminate any excess initial abundance by maintaining equilibrium. In contrast, for freeze-in, due to the weak interaction strength, it is valuable to investigate how such a small initial abundance can extend the duration of the EMD period. Assuming an initial temperature at $T_{i}$, let’s consider a scenario where species $S$ has a non-zero initial abundance, denoted by $\eta_{S}$, which is proportional to the ratio of its energy density $\rho_{S}(T_{i})$ to that of the radiation component $\rho_{R}(T_{i})$ at that temperature. Furthermore, we’ll simplify our analysis by assuming that, due to their extremely weak interactions, these particles become non-relativistic shortly after their production, so that one can redshift the initial abundance to get the abundance at the onset of EMD era,
@@
Figure 1: Schematic timeline of key events: Following inflation, a $(B-L)$ phase transition generates cosmic string networks, which become a dominant source of GWs primarily through loop production. Simultaneously, a real scalar singlet, $S$, acquires a mass dependent on the $(B-L)$ Higgs vacuum expectation value. Starting from $H_{\rm EMD}$ to $\Gamma_{S}$, this scalar dominates the universe’s energy density, after which the universe transitions to standard radiation domination. The long-lived scalar leaves an imprint on the high-amplitude GWs generated by cosmic strings, with the duration ($N_{e}$) and end of matter domination ($\Gamma_{S}$) influencing the spectral shapes, which would otherwise produce a scale-invariant, flat plateau at high frequencies. Upcoming beam dump experiments offer a compelling and complementary approach to search for and constrain such long-lived BSM particles based on their mass ($m_{S}$) and decay width ($\Gamma_{S}$) to the SM states.
@@ VII Results
Freeze-in: $H=10^{12}{\rm GeV}$, $\eta_{S}=10^{-13}$
Parameters $\sin\theta$ $m_{S}(\rm GeV)$ $N_{e}$
BP1 $10^{-5.38}$ $10^{0.13}$ 4.15
@@
Table 1: The 3 benchmark points considered in this paper for the modified freeze-in scenario. In each case, there is a possibility of complementary detection by lab searches and GW observations.
@@
Figure 2: Complementarity of GW and laboratory searches (indicated by colored dots) for Top-left: standard non-relativistic FO with local $(B-L)$ strings with $v_{\rm CS}=10^{14}$ GeV, Top-right: standard non-relativistic FO with global $(B-L)$ strings with $v_{\rm CS}=10^{15}$ GeV, Bottom-left: standard FI with $H_{i}=10^{8}$ GeV and local $(B-L)$ strings with $v_{\rm CS}=10^{14}$ GeV, Bottom-right: standard FI with $H_{i}=10^{8}$ GeV and global $(B-L)$ strings with $v_{\rm CS}=10^{15}$ GeV. The colored shaded regions represent the sensitivity curves of various GW experiments, such as $\mu$Ares (red), LISA (blue), DECIGO (green), and ET (brown), with the stipulation that the first turning point frequency to be detected with SNR $>10$. The gray shaded regions are excluded by CHARM, LHCb, CMS re-interpretation results. The solid lines show the sensitivity of upcoming long-lived particle searches like DarkQuest1, DarkQuest2, MATHUSLA, SHiP, etc.
@@
where $\Omega_{\rm noise}(f)$ is the noise curve of a given experiment and $f_{\rm min}$($f_{\rm max}$) are minimum (maximum) accessible frequency. Interestingly, the forthcoming GW detectors are sensitive enough to pinpoint any such imprints, as shown in all subplots of Fig.2 through color-shaded regions for $\mu$Ares Sesana:2019vho (red), LISA LISA:2017pwj (blue), DECIGO Kawamura:2020pcg (green) and ET Punturo:2010zz (brown) respectively, and thus may provide complementary information about the scalar mass and its interactions. An examination of the subplots reveals a significant difference in the sensitivities of GW detectors to EMD for local vs global symmetry breaking (cosmic strings). This discrepancy arises from the differing amplitudes of the corresponding cosmic string networks at high frequencies, as evident in Fig. 6.
Figure 3: Complementarity of GW and laboratory searches with a non-zero initial $S$- abundance(indicated by colored dots) for Top-left: freeze-in with local $(B-L)$ strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-15}$, Top-right: same as the earlier with global $(B-L)$ strings with $v_{\rm CS}=10^{15}$ GeV, Bottom-left: freeze-in with local $(B-L)$ strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-13}$, Bottom-right: same as the earlier with global $(B-L)$ strings with $v_{\rm CS}=10^{15}$ GeV. As above, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
@@
Figure 4: Complementarity of GW and colliders dedicated for higher mass searches with a significantly larger initial $S$- abundance(indicated by colored dots) for Left: freeze-in with local $(B-L)$ strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-5}$, Right: same as the earlier with global $(B-L)$ strings with $v_{\rm CS}=10^{15}$ GeV. As above, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
@@
Figure 5: Imprints of long-lived scalar $S$ on SGWB from cosmic strings and complementarity with collider searches for the benchmarks BP1 ($\star$”), BP2 $\spadesuit$”), BP3 $\clubsuit$”) (see Table-1) shown in Top-left: for local strings, Top-right: for global strings. As above, for both plots, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
Fig. 5 illustrates three benchmark cases (BP1, BP2, and BP3) detailed in Table 1 and represented by $\star$, $\spadesuit$, and $\clubsuit$, respectively. All benchmark points fall well within the projected sensitivity ranges of upcoming long-lived particle searches like SHiP, MATHUSLA Curtin:2023skh etc. The turning point frequencies for all benchmark points, for both local (with $v_{\rm CS}=10^{14}$ GeV) and global string (with $v_{\rm CS}=10^{15}$ GeV) cases, fall within the sensitivity range of current and planned GW searches. Fig.6 presents the SGWB spectra for both string classes, considering these three benchmarks and summing over a large number of modes, $k=1000$. In principle, if one sums over an infinite number of modes there will be a characteristic $f^{-1/3}$ fall after the break frequency ($f_{\rm brk}$) and a second plateau due to the loops formed at the early radiation epoch, for a finite duration of EMD. Although, a prolonged period of EMD started at a very early time ($T_{\rm dom}\gg T_{\rm dec}$) shifts the second kink ($f_{\rm dom}$) to much higher frequencies inaccessible to planned GW detectors, a short EMD may allow for the detection of this second spectral break. Interestingly, Such detection could provide valuable insights into the strength of EMD ($N_{e}$).
Figure 6: The GW spectral shapes for local strings (left panel) and global strings (right panel) are shown with a period of early matter domination that lasts $N_{e}=$ $4.15$ (Solid), $6.8$ (dotted) and $8.05$ (dashed) respectively. We also mark each benchmarks BP1 ($\star$”), BP2 $\spadesuit$”), BP3 $\clubsuit$”) (see Table-1). The period of matter domination is indicated by a kink and decline in the spectrum, whereas the spectrum is relatively flat for frequencies that correspond to periods of radiation domination. Note that we show benchmarks where one timescale is observable as well as other benchmarks where one can observe both the start and end of matter domination.
@@ VIII Discussion and Conclusion
The entire setup is therefore described by a minimal set of independent parameters: the mass and decay width of the scalar represented as, $m_{S}$ and $\Gamma_{S}$, and the energy fraction of $S$ particles at the end of inflation $\eta_{S}$. Assuming several instances of $U(1)$ symmetry breaking scales caused to vev $\langle\Phi\rangle$ of scalar $\Phi$, which via Higgs-portal like mixing $(\lambda_{\Phi S})$ leads to the generation of $S$ mass, we find a novel complementarity signature for long-lived $S$ in GW searches in LISA, ET, $\mu$Ares etc. We studied the effect of those different parameters on the shape of the GW spectrum with a period of early matter domination caused due to long-lived $S$. We showed that such a non-standard cosmological history leaves imprints on the GW signals, to be measured by upcoming detectors. We exhibited the GW spectral shapes that may be observed in LISA, ET, BBO, DECIGO and u-DECIGO.
Because the production of the $S$ particle is intrinsic to our model and implicitly given by the lifetime, mass and decay branching fraction of $S$ into visible sectors, the interaction strength of the $S$ and SM particles (the Higgs for instance) can be simply extracted from each point in the parameter space. We showed that the long-lived particle $S$ can therefore be searched for experimentally in light-dark sector searches involving intensity, lifetime and beam dump experiments. We explored the parameter space in Fig. 5 and compared our results to the projected limits of laboratory searches such as FASER Feng:2017vli ; FASER:2018eoc ; FASER:2018bac ; FASER:2019aik , DUNE DUNE:2015lol ; Berryman:2019dme , DARKQUEST-2 Batell:2020vqn , MATHUSLA Curtin:2018mvb , PS191 Bernardi:1985ny ; Gorbunov:2021ccu , and SHIP SHiP:2015vad . In particular, we found that an early matter-dominated period with equation-of-state parameter $\omega\approx 0$, with $m_{S}\sim\mathcal{O}(0.5-5)$ GeV, could show detectable signals for DUNE, MATHUSLA, and SHIP. Interestingly, we also showed that this same region of parameter region would yield detectable signals within four years of exposure for GW detectors such as LISA, ET and u-DECIGO, serving as a complementary smoking-gun signature.
@@ IX ACKNOWLEDGEMENT
The work of SD is supported by the National Natural Science Foundation of China (NNSFC) under grant No. 12150610460. GW acknowledge the STFC Consolidated Grant ST/L000296/1 a

### v2

@@ VI Complementarity between GW and laboratory searches
Probing the aforementioned long-lived scalar $S$ in upcoming colliders is very challenging. However, in the early universe, non-negligible processes such as inflaton decay or moduli decay can result in a significant initial abundance of $S$, potentially altering the standard scenario. For freeze-out, a sufficiently high interaction rate would eliminate any excess initial abundance by maintaining equilibrium. In contrast, for freeze-in, due to the weak interaction strength, it is valuable to investigate how such a small initial abundance can extend the duration of the EMD period. Assuming an initial temperature at $T_{i}$, let’s consider a scenario where species $S$ has a non-zero initial abundance, denoted by $\eta_{S}$, which is proportional to the ratio of its energy density $\rho_{S}(T_{i})$ to that of the radiation component $\rho_{R}(T_{i})$ at that temperature. Furthermore, we’ll simplify our analysis by assuming that, due to their extremely weak interactions, these particles are initially relativistic but become non-relativistic as their mass approaches the temperature scale, so that one can redshift the initial abundance to get the abundance at the onset of EMD era,
@@
Figure 1: Schematic timeline of key events: Following inflation, a (BL)(B-L) phase transition generates cosmic string networks during the early radiation-dominated (ERD) era, which become a dominant source of GWs primarily through loop production. Simultaneously, a real scalar singlet, $S$, acquires a mass dependent on the (BL)(B-L) Higgs vacuum expectation value. Starting from $H_{\rm EMD}$ to $\Gamma_{S}$, this scalar dominates the universe’s energy density, after which the universe transitions to standard radiation domination. The long-lived scalar leaves an imprint on the high-amplitude GWs generated by cosmic strings, with the duration ($N_{e}$) and end of matter domination ($\Gamma_{S}$) influencing the spectral shapes, which would otherwise produce a scale-invariant, flat plateau at high frequencies. Upcoming beam dump experiments offer a compelling and complementary approach to search for and constrain such long-lived BSM particles based on their mass ($m_{S}$) and decay width ($\Gamma_{S}$) to the SM states.
@@ VII Results
Freeze-in: $H=10^{12}{\>\rm GeV},\eta_{S}=10^{-13}$
Parameters $\sin\theta$ $m_{S}$ (GeV) $N_{e}$
BP1 $^{-5.38} 100.1$$^{0.13}$ 4.15
@@
Table 1: The three benchmark points considered in this paper for the modified freeze-in scenario. In each case, there is a possibility of complementary detection by lab searches and GW observations.
@@
[…]
@@
$\text{SNR}=\sqrt{t_{\rm obs}\int_{f_{\rm min}}^{f_{\rm max}}df\left(\frac{\Omega_{\rm GW}(f)}{\Omega_{\rm noise}(f)}\right)^{2}}>10,$ (34)
@@
Figure 3: Complementarity of GW and laboratory searches with a non-zero initial $S$- abundance(indicated by colored dots) for Top-left: freeze-in with local (BL)(B-L) strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-15}$, Top-right: same as the earlier with global (BL)(B-L) strings with $v_{\rm CS}=10^{15}$ GeV, Bottom-left: freeze-in with local (BL)(B-L) strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-13}$, Bottom-right: same as the earlier with global (BL)(B-L) strings with $v_{\rm CS}=10^{15}$ GeV. As above, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
@@
Figure 4: Complementarity of GW and colliders dedicated for higher mass searches with a significantly larger initial $S$- abundance(indicated by colored dots) for Left: freeze-in with local (BL)(B-L) strings with $v_{\rm CS}=10^{14}$ GeV, initial Hubble scale $H_{i}=10^{12}$ GeV and $\eta_{S}=10^{-5}$, Right: same as the earlier with global (BL)(B-L) strings with $v_{\rm CS}=10^{15}$ GeV. As above, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
@@
Figure 5: Imprints of long-lived scalar $S$ on SGWB from cosmic strings and complementarity with collider searches for the benchmarks BP1 ($\star$”), BP2 $\spadesuit$”), BP3 $\clubsuit$”) (see Table-1) shown in left: for local strings, right: for global strings. As above, for both plots, the shaded regions correspond to the ability of a GW detector to “see” the end of matter domination.
Fig. 5 illustrates three benchmark cases (BP1, BP2, and BP3) detailed in Table 1 and represented by $\star$, $\spadesuit$, and $\clubsuit$, respectively. All benchmark points fall well within the projected sensitivity ranges of upcoming long-lived particle searches like SHiP, MATHUSLA Curtin:2023skh etc. The turning point frequencies for all benchmark points, for both local (with $v_{\rm CS}=10^{14}$ GeV) and global string (with $v_{\rm CS}=10^{15}$ GeV) cases, fall within the sensitivity range of current and planned GW searches. Fig.6 presents the SGWB spectra for both string classes, considering these three benchmarks and summing over a large number of modes, $k=10^{3}$. In principle, if one sums over an infinite number of modes there will be a characteristic $f^{-1/3}$ fall after the break frequency ($f_{\rm brk}$) and a second plateau due to the loops formed at the early radiation epoch, for a finite duration of EMD. Although, a prolonged period of EMD started at a very early time ($T_{\rm dom}\gg T_{\rm dec}$) shifts the second kink ($f_{\rm dom}$) to much higher frequencies inaccessible to planned GW detectors, a short EMD may allow for the detection of this second spectral break. Interestingly, Such detection could provide valuable insights into the strength of EMD ($N_{e}$).
Figure 6: The GW spectral shapes for local strings (left panel) and global strings (right panel) are shown with a period of early matter domination that lasts $N_{e}=$ $4.15$ (Solid), $6.8$ (dotted) and $8.05$ (dashed) respectively. We also mark each benchmarks BP1 ($\star$”), BP2 $\spadesuit$”), BP3 $\clubsuit$”) (see Table-1). The standard cosmological scenario without an EMD phase is shown as gray solid lines in both plots. In contrast, scenarios with an intermediate EMD phase exhibit a distinctive kink and a suppression in the gravitational wave spectrum, marking the period of matter domination, whereas the spectrum is relatively flat for frequencies that correspond to periods of radiation domination. Note that we show benchmarks where one timescale is observable as well as other benchmarks where one can observe both the start and end of matter domination.
@@ VIII Discussion and Conclusion
[…]
Because the production of the $S$ particle is intrinsic to our model and implicitly given by the lifetime, mass and decay branching fraction of $S$ into visible sectors, the interaction strength of the $S$ and SM particles (the Higgs for instance) can be simply extracted from each point in the parameter space. We showed that the long-lived particle $S$ can therefore be searched experimentally in light-dark sector searches involving intensity, lifetime, and beam dump experiments. We explored the parameter space in Fig. 5 and compared our results to the projected limits of laboratory searches such as FASER Feng:2017vli; FASER:2018eoc; FASER:2018bac; FASER:2019aik, DUNE DUNE:2015lol; Berryman:2019dme, DARKQUEST-2 Batell:2020vqn, MATHUSLA Curtin:2018mvb, PS191 Bernardi:1985ny; Gorbunov:2021ccu, and SHIP SHiP:2015vad. In particular, we found that an early matter-dominated period with equation-of-state parameter $\omega\approx 0$, with $m_{S}\sim\mathcal{O}(0.5-5)$ GeV, could show detectable signals for DUNE, MATHUSLA, and SHIP. Interestingly, we also showed that this specific parameter region would yield detectable signals within four years of exposure for GW detectors such as LISA, ET and u-DECIGO, serving as a complementary smoking-gun signature.
@@ IX ACKNOWLEDGEMENT
The work of SD is supported by the National Natural Science Foundation of China (NNSFC) under grant No. 12150610460. GW acknowledge the STFC Consolidated Grant ST/L000296/1 a.